To attract more visitors, zoo authority announces 20% discount on every ticket which cost 29 paise. For this reason sale increases by 28% . Find the percentage of increase in the number of visitors.


1
Expert's answer
2021-02-16T09:47:23-0500

Let the number of visitors in beginning = 100 (say)

"\Rightarrow" Total revenue = 0.29 "\times" 100 = 29 Rupees

New price after discount = 29 "\times" "\dfrac{80}{100}"

"\rightarrow" 23.2 Paise or 0.232 Rupees

New revenue = 29 "\times" "\dfrac{128}{100}"

"\rightarrow" 37.12 Rupees

Number of visitors now = "\dfrac{37.12}{0.232}" = 160

% change in number of visitors ="\dfrac{160-100}{100}" = 60


Hence, 60% of increase in number of visitors.

"\dfrac{}{}"